zarkam21:

Questions 8 - 10, deal with the ambiguous SSA case. For each, find all possible solutions and sketch the triangle(s) in each case.

zarkam21:

1 attachment
Vocaloid:

for SSA we have two possibilities depending on the arrangement of the side across from the angle

Vocaloid:

so first you'd use law of sines to find angle B using the left triangle, solve for the remaining sides/angles

Vocaloid:

as a reminder: a/sin(A) = b/sin(B) for law of sines

zarkam21:

15/sin(50) = 12/sin(B)

Vocaloid:

awesome, now solve for B

zarkam21:

0.66 2.48

zarkam21:

DO i round to the nearest tenth or hundredth?

Vocaloid:

you should only be getting 1 value for B also the solution should be in degrees not radians since the angle you are given is also in degrees try cross multiplying 15/sin(50) = 12/sin(B) and solving for sin(B), then take the arcsin of the result

zarkam21:

4 sin (50)/5

Vocaloid:

good, then take the arcsin of 4 sin (50)/5

zarkam21:

0.61

zarkam21:

i think :/

zarkam21:

wait 39.79

Vocaloid:

remember your answer should be in degrees not radians

Vocaloid:

good, 37.79 degrees for angle B now you have angle A and angle B, find angle C

zarkam21:

i use the c^2=a^2+b^2-2ab cosC right

Vocaloid:

hint: all the triangles add up to 180 degrees you can't use law of cosines yet because you don't have c

Vocaloid:

*all the angles in a triangle

zarkam21:

Im not sure i I would use the sin (a)-a formula or just subtract 180-

zarkam21:

a/sin(A) = b/sin(B)

Vocaloid:

angle A + angle B + angle C = 180 so angle C = ?

zarkam21:

something like this right

zarkam21:

oh okay 180-37.79-50=92.21

zarkam21:

for angle c

Vocaloid:

good now you have angle C, you can now use law of cosines to find c

zarkam21:

okay c^2=a^2+b^2-2ab (cos(C)) c^2=15^2+12^2-2*15*12(cos(C)

zarkam21:

like this

Vocaloid:

good but you need to plug in angle C = 92.21

zarkam21:

oh right..

zarkam21:

c^2=15^2+12^2-2*15*12(cos(92.21) c^2=531.07 c=24.05

Vocaloid:

wait a min

zarkam21:

sure.

Vocaloid:

awesome, so that's the first ambiguous case solved however, for the SSA case, angle B can also be obtuse [see the diagram] so we take 180 - the old angle B to get the new angle B so 180 - 92.21 = 87.79 as the new angle C then we re-solve the triangle for angle B and side c

Vocaloid:

so if angle A = 50 and angle C = 87.79 what is the new angle B?

zarkam21:

we use law of cosines right

zarkam21:

wait nvm its 180-50-87.79=42.21

Vocaloid:

yeah i just double checked and apparently there's only one case for this triangle so it's just the other set of solutions (angle B = 37.79, angle C = 92.2, c = 24.05)

Vocaloid:

so yeah that's #8 done, repeat the process for #9 and start by finding angle B using law of sines

zarkam21:

okay so is just the orgiinal ones we found right

zarkam21:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid awesome, so that's the first ambiguous case solved however, for the SSA case, angle B can also be obtuse [see the diagram] so we take 180 - the old angle B to get the new angle B so 180 - 92.21 = 87.79 as the new angle C then we re-solve the triangle for angle B and side c \(\color{#0cbb34}{\text{End of Quote}}\)

zarkam21:

not this

Vocaloid:

yeah not that

zarkam21:

a/sin(A) = b/sin(B) 15/sin(50) = 12/sin(B) 37.79 degrees for angle B 180-37.79-50=92.21 (Angle C) c^2=15^2+12^2-2*15*12(cos(92.21) c^2=531.07 c=24.05

zarkam21:

so just this

zarkam21:

okay we can continue now

Vocaloid:

yes

zarkam21:

a/sin(A) = b/sin(B) 10/sin(50) = 15/sin(B)

zarkam21:

3sin(50)/2

zarkam21:

\[\frac{ 3 }{ 2 } \cos(\frac{ 2\pi }{ 9})\]

zarkam21:

72.64 degrees

Vocaloid:

so you got sin(B) = 3sin(50)/2 you just need to take the arcsin of both sides to find what B is, converting to cos isn't helpful here

zarkam21:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zarkam21 72.64 degrees \(\color{#0cbb34}{\text{End of Quote}}\)

zarkam21:

is this it

Vocaloid:

huh for some reason it's giving me B = 90 degrees

zarkam21:

its prob me thats wrong

Vocaloid:

oh duh since 3sin(50)/2 is greater than 1 then there's no solution for the arcsin making this triangle have no real solutions

zarkam21:

a/sin(A) = b/sin(B) 10/sin(50) = 15/sin(B) 72.64 degrees for angle B 180-72.64-50=57.36(Angle C) c^2=10^2+15^2-2*10*15(cos(57.36) c^2=118.44 c=10.88

zarkam21:

so this is all not necessary?

Vocaloid:

yeah, angle B has no solutions meaning this triangle cannot actually exist moving on to #10 i guess, we'd solve for angle B as usual using law of sines

zarkam21:

a/sin(A) = b/sin(B) 1.6/sin(50) = 2/sin(B) =-6.1

zarkam21:

6.28

Vocaloid:

the angle is 47 not 50

Vocaloid:

arcsin(2sin(47)/1.6) = ?

zarkam21:

1.15

Vocaloid:

good but what is that in degrees not radians?

zarkam21:

69.89

Vocaloid:

a little off, 66.09 degrees now find angle C, given angle A = 47 and angle B = 66.09

zarkam21:

180-66.09-47=66.91 (Angle C) c^2=1.6^2+2^2-2*1.6*2(cos(66.91) c^2=10.35 c=3.22

Vocaloid:

yeah that's good

Vocaloid:

I was double checking w/ law of sines and for some reason I get a diff. value for c hm..

zarkam21:

oh really

zarkam21:

so is the final answer okay or just i recheck

zarkam21:

should*

Vocaloid:

ok apparently you can't use the law of cosines for an ambigious case, in that case we would simply use law of sines to find c c/sin(C) = a/sin(A) so c/sin(66.91) = 1.6/sin(47) gives the value of c

Vocaloid:

that means we need to re-do the value of c for triangle #8 but we can get back to that after we finish #10 i guess

Vocaloid:

c/sin(66.91) = 1.6/sin(47) multiplying both sides by sin(66.91) gives us c = 1.6 * sin(66.91) / sin(47) = ?

zarkam21:

2.01

zarkam21:

this would be #10 right

Vocaloid:

good but for this triangle we have another possible solution first we take 180 - angle B to get 180 - 66.09 = 113.91 = the other possible angle B value so for this new case, we have angle A = 47, angle B = 113.91, angle C = ?

zarkam21:

180-113.91-47=19.09

Vocaloid:

awesome, then use law of sines again to find c c/sinC = a/sinA, plug in angles C and A, plus side a to solve for side c

zarkam21:

c/sinC = a/sinA c/sin(19.09)/1.6/sin(47)

Vocaloid:

that / in the middle should be an equal sign c/sin(19.09) = 1.6/sin(47) now find c

zarkam21:

2.19

zarkam21:

0.72

zarkam21:

*

Vocaloid:

good, 0.72 = c giving us the other solution for the ambiguous case anyway, going back to #8, we found that angle C = 92.21 giving us c/sin(92.21) = 15/sin(50) giving us c = 19.57 anyway, let's do a quick recap of our solutions:

Vocaloid:

#8: only one solution angle B = 37.79 degrees angle C = 92.21 degrees, c = 19.57 #9: no sol'ns (the arcsin of angle B is greater than 1 so no possible angle B values) #10: two solutions solution 1: angle B = 66.09 degrees, angle C = 66.91 degrees, c = 2.01 solution 2: angle B = 113.91 degrees, angle C = 19.09 degrees, c = 0.72

Vocaloid:

and that's it

zarkam21:

oh for #8, we got c^2=531.07 c=24.05

zarkam21:

its 19.57 instead?

Vocaloid:

yeah, I realized we can't use law of cosines we have to use law of sines which gives us 19.57